### v1

# Co-Seg++: Mutual Prompt-Guided Collaborative Learning for Versatile Medical Segmentation

Abstract
Medical image analysis is critical yet challenged by the need of jointly segmenting organs or tissues, and numerous instances for anatomical structures and tumor microenvironment analysis. Existing studies typically formulated different segmentation tasks in isolation, which overlooks the fundamental interdependencies between these tasks, leading to suboptimal segmentation performance and insufficient medical image understanding. To address this issue, we propose a Co-Seg++ framework for versatile medical segmentation. Specifically, we introduce a novel co-segmentation paradigm, allowing semantic and instance segmentation tasks to mutually enhance each other. We first devise a spatio-temporal prompt encoder (STP-Encoder) to capture long-range spatial and temporal relationships between segmentation regions and image embeddings as prior spatial constraints. Moreover, we devise a multi-task collaborative decoder (MTC-Decoder) that leverages cross-guidance to strengthen the contextual consistency of both tasks, jointly computing semantic and instance segmentation masks. Extensive experiments on diverse CT and histopathology datasets demonstrate that the proposed Co-Seg++ outperforms state-of-the-arts in the semantic, instance, and panoptic segmentation of dental anatomical structures, histopathology tissues, and nuclei instances. The source code is available at https://github.com/xq141839/Co-Seg-Plus.
{IEEEkeywords}
Collaborative Learning, medical semantic segmentation, medical instance segmentation, mutual prompt
1 Introduction
[…]
Medical image segmentation plays a crucial role in clinical applications and has received significant research attention. [1, 2, 3]. For example, histopathology image analysis demands accurate segmentation of tissue regions and separation of individual nuclei within these regions to assess tissue subtypes and tumor grading [4]. Similarly, cone beam computed tomography (CBCT) requires precise delineation of anatomical structures such as jaw bones and sinuses, alongside instance-level segmentation of individual teeth for dental treatment planning [5]. These dual requirements have given rise to the challenging field of versatile medical segmentation, which seeks to unify both semantic and instance segmentation tasks across various medical imaging contexts.
Figure 1: Comparison of our Co-Seg++ and existing segmentation works. (a) Two independent networks for semantic and instance segmentation. (b) A shared image encoder but separated task decoders for semantic and instance segmentation. (c) Our Co-Seg++ leverages spatio-temporal prompts for collaborative semantic and instance segmentation.
Figure 2: Regional interdependencies between semantic and instance segmentation tasks in nuclei images. Semantic regions (e.g., glands) often encompass multiple types of instance-level structures (e.g., lympho-reticular (LR) and connective (Con) nuclei). The relationship of semantic and instance segmentation confirms contextual understanding and motivates the proposed joint optimization of versatile medical segmentation.
Typically, medical image segmentation frameworks [6, 7, 8] have been designed for specific tasks, resulting in isolated solutions that address semantic segmentation and instance segmentation separately, as shown in Fig. 1(a). These conventional approaches typically adopt independent encoder-decoder architectures without shared information during the training process. The semantic segmentation networks are optimized to classify each pixel into pre-defined categories, meanwhile instance segmentation models [9, 10, 2] focus on detecting and delineating individual objects. However, multiple independent networks [11, 12] lead to increased computation complexity. More seriously, this separation learning [13] impairs consistency for similar low-level features across tasks, resulting in redundant feature extraction. These limitations motivate the development of more integrated approaches that could leverage commonalities between semantic and instance segmentation tasks in medical images.
To this end, the encoder-shared framework has emerged as a promising alternative paradigm for versatile medical segmentation, as shown in Fig. 1(b). These methods [14, 15] utilize a common feature extraction backbone for both semantic and instance segmentation tasks, enabling more efficient resource utilization and establishing a foundation for potential information sharing. The shared encoder captures universal representations beneficial to both tasks, reducing computational redundancy and providing consistency in the initial feature extraction stage. While this paradigm improves efficiency compared to fully independent models, it still maintains separate decoders, leading to limited cross-task interaction during the critical decoding phase, which fails to exploit the mutual guidance that could be derived from intermediate representations of each task. In fact, we observe that both segmentation tasks are highly correlated, by analyzing the regional interdependencies of multiple histopathology datasets [16, 17, 14] in Fig. 2. These segmentation tasks aim to achieve an adequate perception and understanding of histopathology images. For example, accurately identifying nuclei provides valuable cues for understanding the underlying tissue structures, whereas tissue segmentation can aid in localizing nuclei. These synergies motivate us to develop a collaborative framework that jointly optimizes semantic and instance segmentation to improve medical image analysis.
To overcome this bottleneck, we propose Co-Seg++, a collaborative versatile medical segmentation framework that allows semantic and instance segmentation tasks to mutually enhance each other. As illustrated in Fig. 1(c), Co-Seg++, based on a novel co-segmentation paradigm, improves segmentation mask quality by capturing contextual dependencies between the two tasks. Specifically, we design a spatio-temporal prompt encoder (STP-Encoder) that adaptively integrates spatial context and temporal dynamics to generate prior spatial constraints for guiding segmentation decoding across both tasks. We further introduce a multi-task collaborative decoder (MTC-Decoder) that leverages cross-task guidance to jointly predict semantic and instance segmentation masks in a unified framework. This collaborative learning strategy not only achieves enhanced contextual consistency but also significantly reduces prediction errors through mutual task optimization. Extensive experiments on diverse histopathology and dental datasets demonstrate that our Co-Seg++ consistently outperforms state-of-the-art segmentation methods.
The contributions of this work are summarized as follows:
  • We propose Co-Seg++, a versatile medical segmentation framework that incorporates a novel co-segmentation paradigm, enabling semantic and instance segmentation tasks to mutually enhance each other.
  • We devise an STP-Encoder that integrates long-range spatial and temporal information to extract high-quality semantic and instance prompts, providing effective prior knowledge for cross-task segmentation guidance.
  • We design a MTC-Decoder that employs shared image embeddings and joint prompt mechanisms for collaborative semantic and instance segmentation decoding, achieving mutual task optimization and enhanced mask prediction consistency.
  • We conduct extensive experiments on diverse medical image datasets, and our Co-Seg++ framework outperforms state-of-the-art independent and encoder-shared medical segmentation methods across semantic, instance, and panoptic segmentation tasks.
Figure 3: The overview of our Co-Seg++ framework for collaborative versatile medical segmentation. For ease of understanding, we elaborate on the case of Co-Seg++ in tissue and nuclei segmentation. Our Co-Seg++ fully exploits complementary context by leveraging spatio-temporal prompts.
2 Related Work
2.1 Medical Image Segmentation
Medical image segmentation is fundamental to computer-aided diagnosis and treatment planning [6, 7, 8]. This field encompasses semantic segmentation (i.e., pixel-level classification), instance segmentation (i.e., individual object delineation), and panoptic segmentation that unifies both by jointly predicting semantic classes and instance identities [18]. Building on the U-Net encoder-decoder framework [6], recent advances span three network architectures: convolutional neural network (CNN), vision transformers (ViT) [19], and more recent Mamba models [20]. The CNN-based methods [7, 21, 8, 22] applied automated architecture optimization and multi-scale convolutions to enhance target awareness. The ViT-based methods [23, 24] leveraged self-attention mechanisms to capture global contexts. The Mamba-based methods [12, 25, 26] utilized sequence modeling to handle long-range relationships with linear computational complexity. These works have demonstrated strong adaptability across diverse imaging modalities.
Histopathology and CBCT represent key modalities for tumor assessment and dental practices, with applications spanning semantic and instance segmentation. In histopathology, semantic segmentation localizes tissue regions while instance segmentation identifies individual nuclei for tumor microenvironment profiling [9, 27, 11]. Methods like CPP-Net [10] used multi-point sampling for polygon distance mapping, and CellViT [2] employed multi-task learning for multi-directional distance maps. For CBCT analysis, semantic segmentation parses orofacial structures while instance segmentation recognizes individual teeth [28]. PadUNet [29] achieved 3D nerve canal segmentation via sparse annotations, while TSG-GCN [30] optimized spatial tooth distribution through dynamic adjacency learning. However, existing approaches employ task-specific models. Our Co-Seg++ transcends these limitations, achieving versatile segmentation with superior performance.
2.2 Versatile Segmentation Paradigms
The classical versatile medical segmentation frameworks have predominantly adopted a task-decoupled paradigm to address multiple segmentation objectives simultaneously. The TIAToolbox [13] employed an ensemble approach that integrates multiple independent networks for tissue and nuclei segmentation in histopathology images, treating each task as a separate optimization problem. To enhance cross-task consistency, Cerberus [14] employed the shared encoder with decoupled decoders for joint training efficiency. Federico et al. [31] adopted multi-class strategies with independent labels for CBCT tasks. However, they lack flexible guidance mechanisms that limit their ability to incorporate contextual knowledge or interactive signals, highlighting the need for more adaptive and guidance-driven segmentation paradigms.
The segment anything model (SAM) [32] introduced prompt-based segmentation through point, box, text, or mask guidance. Recent studies [3, 33, 34, 35, 36] have demonstrated the strong adaptability of SAM across a variety of medical scenarios. These prompt-driven methods enabled flexible segmentation by incorporating domain knowledge directly into the inference process. To further refine segmentation outputs, PromptNucSeg [1] utilized an anchor-based detection prompter to automatically produce multi-point prompts within each nucleus for deeply guiding nuclei segmentation. UN-SAM [37] devised a multi-scale self-prompting strategy to extract coarse mask prompts from image embeddings, eliminating the demand for labor-intensive manual annotations. However, these prompt-based approaches are designed to improve performance within individual segmentation tasks. Unlike existing approaches that utilize prompts for task-specific enhancement, Co-Seg++ explores complementarities between semantic and instance segmentation in medical scenarios, which enables both tasks to mutually guide each other.
3 Methodology
We present the Co-Seg++ framework with the co-segmentation paradigm in Fig. 3, and achieve the mutual optimization between semantic and instance segmentation for medical images. To accomplish this, we design the STP-Encoder to provide prior spatial constraints to guide both tasks, and the MTC-Decoder to cooperatively generate semantic and instance maps through bidirectional information interactions, which enables the two tasks to enhance each other, leveraging spatio-tempral prompts for improved segmentation accuracy.
Figure 4: The illustration of STP-Encoder, integrating spatial and temporal prompts to establish long-range relationships between target segmentation regions and shared image embeddings, providing prior spatial constraints for cross-task guidance.
3.1 Co-Segmentation Paradigm
Existing versatile medical segmentation methods [14, 15] decouple the parameter space of semantic and instance segmentation, disrupting their interdependencies, degrading the performance. To address this issue, we propose the co-segmentation paradigm that leverages the closed-loop bidirectional interaction to realize dual-task collaborative optimization. Specifically, our goal is to train a model $f_{\Theta}:x\rightarrow\{y_{1},y_{2}\}$, where $\Theta=\{\theta_{1},\theta_{2}\}$ represents the parameters for semantic segmentation $\theta_{1}$ and instance segmentation $\theta_{2}$, with $y_{1}$ and $y_{2}$ as the corresponding segmentation masks. To model the interdependence between the two tasks, we define their joint prediction of $y_{1}$ and $y_{2}$ as follows:
$\begin{split}p(y_{1},y_{2}|x,\theta_{1},\theta_{2})&=p(y_{1}|x,\theta_{1})\cdot p(y_{2}|y_{1},x,\theta_{2})\\ &=p(y_{2}|x,\theta_{2})\cdot p(y_{1}|y_{2},x,\theta_{1}).\end{split}$ (1)
Note that Eq. (1) reveals the mutual dependency between semantic and instance segmentation through intertwined conditional probabilities. It states that $p(y_{1},y_{2}|x,\theta_{1},\theta_{2})$, the joint probability of obtaining both segmentations given the image and model parameters, decomposed in two symmetrical ways:
  • $p(y_{1}|x,\theta_{1})\cdot p(y_{2}|y_{1},x,\theta_{2})$ suggests that once the semantic segmentation $y_{1}$ is known, it directly influences the prediction of the instance segmentation $y_{2}$.
  • Similarly, $p(y_{2}|x,\theta_{2})\cdot p(y_{1}|y_{2},x,\theta_{1})$ implies that knowing the instance segmentation $y_{2}$ affects the outcome of the semantic segmentation $y_{1}$.
Both expressions are mathematically equivalent because they describe the same joint probability from different perspectives, highlighting that each segmentation task provides crucial context that enhances the accuracy of the other. This reciprocal relationship underscores the need for a collaborative approach, where both segmentation tasks are optimized together rather than in isolation, leveraging the full context available from each task to improve overall segmentation results. The gradient calculations of the relationship for $\theta_{1}$ and $\theta_{2}$ is defined as:
$\begin{split}\nabla_{\theta_{i}}\ell_{i}=&-\nabla_{\theta_{i}}\mathbb{E}\left[\log p(y_{i}|x,\theta_{i})\right]\\ &-\nabla_{\theta_{i}}\mathbb{E}\left[\log p(y_{j}|y_{i},x,\theta_{j})\right]\\ &+\nabla_{\theta_{i}}I(y_{1};y_{2}|x),\end{split}$ (2)
where $i,j\in\{1,2\},i\neq j$ and $\nabla_{\theta}I(y_{1};y_{2}|x)=\mathbb{E}\left[\nabla_{\theta}\log\frac{p(y_{1},y_{2}|x)}{p(y_{1}|x)p(y_{2}|x)}\right]$. This term captures the mutual information between $y_{1}$ and $y_{2}$, ensuring that both segmentation tasks benefit from shared feature learning. The first term of Eq. (2) represents the main segmentation loss gradient, while the second term incorporates the implicit gradient of task interdependencies, modeling the complementarity between $\theta_{1}$ and $\theta_{2}$. Note that although $p(y_{j}|y_{i},x,\theta_{j})$ is parameterized by $\theta_{j}$, we include its gradient with respect to $\theta_{i}$ to reflect the implicit dependency introduced by co-training and shared feature representations during optimization. As such, this reflects cross-task interactions arising from shared components (e.g., shared encoder or joint decoder), enabling gradient flow between tasks. The third term captures the mutual information between the two outputs, ensuring both tasks benefit from shared learning signals. This leads to the following optimization update rule:
$\theta_{1}\leftarrow\theta_{1}-\eta\nabla_{\theta_{1}}\ell_{1},\quad\theta_{2}\leftarrow\theta_{2}-\eta\nabla_{\theta_{2}}\ell_{2},$ (3)
where $\eta$ is the learning rate. In this way, the co-segmentation paradigm breaks the barrier of isolating gradient flows from each other in multi-task learning, enabling our Co-Seg++ framework to achieve the mutual optimization of semantic and instance segmentation.
Figure 5: The architecture of the MTC-Decoder, employing cross-guidance mechanisms and probability distribution alignment to enable semantic and instance segmentation tasks to mutually enhance each other while ensuring spatial consistency in segmentation decoding.
3.2 Spatio-Temporal Prompt Encoder
The existing prompt-based segmentation methods [32, 38] typically employ two distinct types of guidance signals: the sparse prompts (e.g., point and box) and the dense prompts (e.g., masks). While the sparse prompts provide precise region-of-interest cues, they offer limited contextual information for complex segmentation tasks compared to dense prompts. The dense prompts, in contrast, provide more comprehensive spatial cues but typically lack temporal context or task-specific structure. The proposed co-segmentation paradigm requires establishing implicit interdependencies between semantic and instance segmentation tasks. To achieve this, we introduce the STP-Encoder that leverages region masks and enriched image features to generate prior spatial constraints, effectively incorporating spatial and temporal prompts to improve the ability of the Co-Seg++ to coordinate between semantic and instance segmentation tasks.
As elaborated in Fig. 4, STP-Encoder comprises two parallel branches designed to capture temporal $t$ and task-specific spatial $s_{i}$ prompts. The temporal branch adopts shared image embeddings $h$ as input and first applies a linear layer $\rm Linear_{down}$ for down-projection. Then, a $3\times 3$ 1D convolution followed by a state space model (SSM) extracts temporal features and models long-range relationships. Finally, we apply another linear layer $\rm Linear_{up}$ for up-projection. These operations can be summarized as follows:
$t={\rm Linear_{up}(LN(SSM(Conv1D(Linear_{down}}(h))))),$ (4)
where $\rm LN(\cdot)$ denotes LayerNorm. Different from the existing Mamba networks [12, 24, 20], this temporal branch omits SiLU activation and gating operations since the image embeddings have already been optimally processed by the image encoder, avoiding redundant feature filtering. Moreover, the spatial branch employs several 2D convolutions to transform task-specific mask logits $m_{i}$ to patch embeddings, followed by self-attention $\rm SA(\cdot)$ to capture global context:
$s_{i}={\rm SA(Conv2D_{2}(GELU(LN(Conv2D_{1}}(m_{i}))))),$ (5)
[…]
$c_{i}={\rm CrossAttention}(Q(s_{i}),K(t),V(t)).$ (6)
On this basis, we apply STP-Encoder to semantic and instance mask logits, effectively obtaining prior spatial constraints $c_{i}$ of Co-Seg++ for guiding versatile segmentation decoding.
3.3 multi-task collaborative decoder
Following the co-segmentation paradigm, we propose the MTC-Decoder to enforce bidirectional interactions between semantic and instance segmentation, ensuring mutual refinement rather than treating them as isolated tasks, as illustrated in Fig. 5. Specifically, the MTC-Decoder contains semantic and instance heads that leverage two sets of query embeddings: $q_{i}$ and $q_{j}$ to save the decoding information of both tasks. To leverage complementary effects between tasks, we first perform self-attention on each query, followed by cross-attention with prior spatial constraints $c_{j}$, as follows:
$q^{\prime}_{i}={\rm CrossAttention}(Q({\rm SA}(q_{i})),K(h\oplus c_{j}),V(h\oplus c_{j})),$ (7)
where $\oplus$ represents the element-wise addition. After that, we apply an MLP to refine the query embeddings on $q^{\prime}_{i}$, and further conduct the reverse cross-attention to generate task-specific image embeddings $g$, as follows:
$g_{i}={\rm CrossAttention}(Q(h\oplus c_{j}),K(q^{\prime}_{i}),V(q^{\prime}_{i})).$ (8)
In this way, the prior spatial constraints $c_{j}$ of instances can improve the boundary awareness of semantic segmentation, while semantic feedback provides region constraints for instance discrimination. To further enhance spatial consistency between complementary tasks, we transform $g_{i}$ to a category probability distribution $y_{i}^{\rm prob}$ using $1\times 1$ convolution and softmax operations. We adopt Kullback-Leibler divergence to compute spatial consistency constraint loss $\mathcal{L}_{\rm SCC}$, as follows:
$\mathcal{L}_{\rm SCC}=\mathbb{E}_{x\sim y_{i}^{\rm prob}(x)}\left[\log_{2}y_{i}^{\rm prob}(x)-\log_{2}y_{j}^{\rm prob}(x)\right].$ (9)
Following the standard SAM [32], each segmentation head adopts a pixel decoder to upsample the refined task-specific image embedding. Finally, our MTC-Decoder generates semantic and instance segmentation predictions by performing the dot product between their unscaled image embeddings and corresponding task queries. The predictions of MTC-Decoder contain both the low-resolution binary mask $y^{\rm bin}$ and the final segmentation mask, where $y^{\rm bin}$ is supervised by binary cross-entropy loss as follows:
$\mathcal{L}_{\rm BM}=-y^{\rm bin}_{i}\log(\hat{y}_{i}^{\rm bin})-y^{\rm bin}_{j}\log(\hat{y}_{j}^{\rm bin}).$ (10)
To optimize the final segmentation masks, we utilize the combination of cross-entropy loss and Dice loss for semantic segmentation and follow [2] for instance segmentation:
$\begin{split}&\mathcal{L}_{\rm SEG}^{\rm sem}=\mathcal{L}_{\rm Dice}+\mathcal{L}_{\rm CE},\\ &\mathcal{L}_{\rm SEG}^{\rm ins}=\mathcal{L}_{\rm Dice}+\mathcal{L}_{\rm Focal}+\mathcal{L}_{\rm MSE}+\mathcal{L}_{\rm MSGE}.\end{split}$ (11)
In this way, the proposed MTC-Decoder adopts the bidirectional interactions of semantic and instance segmentation information to mutually improve the quality of segmentation masks of our Co-Seg++ framework.
Table 1: Comparison with state-of-the-arts on histopathology semantic segmentation.
PUMA CRAG GlaS
Methods Tumor Stroma Gland Gland
Dice mIoU HD Dice mIoU HD Dice mIoU HD Dice mIoU HD
nnUnet [7] 91.18 85.39 244.71 53.45 46.43 244.71 88.61 85.42 184.71 89.71 82.81 148.20
Swin-Umamba [12] 88.30 81.04 272.98 53.78 42.14 336.21 86.92 82.58 244.62 88.49 80.99 146.84
EMCAD [8] 87.68 80.07 286.11 52.70 44.92 343.97 87.84 83.68 252.69 85.36 77.50 210.11
H-SAM [3] 91.31 85.47 249.35 54.30 47.14 370.92 88.92 85.77 222.30 86.84 79.87 193.29
Zig-RiR [39] 90.91 85.20 241.56 55.34 48.32 242.71 88.35 84.94 199.28 86.03 78.83 212.74
HoverNet [9] 86.65 78.72 346.51 46.99 39.13 410.99 84.67 79.41 257.99 84.12 75.31 264.28
HookNet [40] 86.84 79.03 321.59 47.25 39.44 395.72 84.89 79.68 245.83 84.35 75.58 251.94
CellPose [11] 85.92 77.58 358.42 45.73 37.89 425.67 83.74 78.12 271.35 83.21 74.02 278.59
CPPNet [10] 87.34 79.78 312.87 48.56 40.92 378.24 85.43 80.34 231.76 85.67 76.89 234.15
CellViT [2] 90.34 84.12 261.78 53.67 46.89 318.45 88.95 85.23 195.67 89.48 82.74 162.35
PromptNucSeg [1] 91.57 86.08 228.92 55.21 48.73 285.63 89.67 86.49 178.24 89.85 83.07 159.89
PathoSAM [15] 91.78 86.34 215.31 54.86 48.15 292.17 89.28 85.94 183.45 89.93 83.68 156.72
Co-Seg++ 93.15 88.03 193.34 60.88 53.65 231.56 90.78 87.94 172.85 91.56 85.47 135.28
Table 2: Comparison with state-of-the-arts on CBCT semantic segmentation.
Methods Lower Jawbone Upper Jawbone Pharynx
Dice mIoU HD Dice mIoU HD Dice mIoU HD
nnUnet [7] 94.55 91.37 54.64 89.28 73.98 81.31 97.00 94.32 22.96
Swin-Umamba [12] 93.59 90.31 53.07 88.92 86.99 70.64 97.01 94.34 27.30
EMCAD [8] 94.21 90.70 90.84 87.95 85.85 69.57 96.87 94.09 23.12
H-SAM [3] 94.52 91.20 61.67 89.61 87.66 96.25 96.54 93.47 28.16
Zig-RiR [39] 94.26 91.05 60.76 89.70 87.71 80.07 96.83 93.81 15.33
ToothNet [28] 92.73 88.96 67.89 86.54 82.71 89.42 96.28 93.58 32.71
HMG-Net [41] 93.21 89.47 63.25 87.19 83.94 85.78 96.61 93.97 31.89
SGA-Net [42] 93.48 89.82 61.14 87.92 84.58 82.35 96.61 94.79 29.38
T-Mamba [24] 94.67 91.58 48.25 90.12 88.19 65.84 97.05 94.63 21.25
ToothSC-SAM [36] 94.78 91.84 45.78 90.47 88.65 62.15 97.14 94.66 19.09
TSG-GCN [30] 94.65 91.72 47.21 90.23 88.42 64.38 97.08 94.71 20.15
Co-Seg++ 96.54 93.85 38.67 92.78 90.94 52.45 97.73 95.58 13.27
3.4 Optimization Pipeline
To construct our Co-Seg++ framework, we first adopt Hirea ViT [43] as the shared image encoder for versatile medical segmentation, ensuring the consistency of feature learning. In particular, we load [38] to initialize corresponding parameters and freeze these weights to preserve pre-trained knowledge. Additionally, we insert Adapter [44] into FFN layers to achieve parameter-efficient fine-tuning in medical domains. Based on the co-segmentation paradigm, Co-Seg++ includes two decoding forwards, where the semantic and instance heads of the MTC-Decoder share parameters in these two stages. For the first forward, MTC-Decoder leverages the image embedding generated from the image encoder to calculate binary segmentation masks: $y^{\rm bin}_{i}$ and $y^{\rm bin}_{j}$ of both tasks without any prompts. Then, we deliver them to the STP-Encoder to provide prior spatial constraints: $c_{i}$ and $c_{j}$ for both tasks. In the second forward, they will be directly transferred back to the MTC-Decoder to co-generate semantic and instance segmentation masks through cross-guidance. The training of Co-Seg++ consists of two parts: (1) the optimization of probability distribution and prior spatial constraints (2) the joint optimization of semantic and instance segmentation. The overall loss is formulated as:
$\mathcal{L}_{\rm CoSeg}=\underbrace{\lambda_{1}(\mathcal{L}_{\rm SCC}+\mathcal{L}_{\rm BM})}_{\text{Forward 1}}+\underbrace{\mathcal{L}_{\rm SEG}^{\rm sem}+\mathcal{L}_{\rm SEG}^{\rm ins}}_{\text{Forward 2}},$ (12)
where $\lambda_{1}$ are factors to adjust the contribution of the first forward. By optimizing $\mathcal{L}_{\rm CoSeg}$, our framework enables accurate and consistent semantic and instance segmentation through joint learning with spatio-temporal priors and cross-task guidance. Co-Seg++ maintains computational efficiency by leveraging a shared encoder, reusing features across tasks, and introducing only lightweight adapter modules for medical adaptation. This collaborative design not only avoids redundant computation but also allows the two tasks to mutually reinforce each other, leading to more robust predictions. We now proceed to demonstrate the effectiveness of Co-Seg++ through comprehensive experiments across diverse datasets.
Table 3: Comparison with state-of-the-arts on histopathology instance segmentation.
PUMA CRAG GlaS
Methods Tumor Lymph Other Cells Epithelial LR Connective Epithelial LR Connective
F1 AJI F1 AJI F1 AJI F1 AJI F1 AJI F1 AJI F1 AJI F1 AJI F1 AJI
nnUnet [7] 65.84 52.17 54.38 33.62 19.76 12.43 58.92 46.25 43.85 31.74 56.73 45.82 46.52 32.89 55.38 35.62 24.17 15.83
Swin-Umamba [12] 68.73 56.42 57.29 36.85 22.64 14.76 61.48 49.33 46.52 34.17 59.35 48.27 49.16 35.44 57.93 38.29 26.75 17.92
EMCAD [8] 70.15 58.93 58.74 38.42 24.83 16.29 63.27 51.08 48.16 35.73 61.29 50.14 51.38 37.62 59.74 40.16 28.49 19.35
H-SAM [3] 74.32 62.58 61.47 40.93 28.56 18.74 67.85 55.29 52.84 39.16 65.73 54.37 56.29 41.85 62.47 42.93 33.18 22.67
Zig-RiR [39] 71.89 60.24 59.83 39.57 26.47 17.35 65.18 52.76 50.32 37.48 63.41 52.15 53.74 39.17 60.83 41.52 30.64 20.81
HoverNet [9] 74.86 63.42 62.35 41.18 28.67 19.15 68.73 55.94 53.81 40.26 66.94 55.47 57.92 42.74 64.29 43.85 34.18 23.59
HookNet [40] 72.61 60.63 61.21 39.80 24.12 15.28 64.74 52.16 49.79 36.29 62.27 50.98 51.69 37.33 61.16 39.77 29.83 19.68
CellPose [11] 75.34 63.91 63.17 41.73 29.45 19.82 69.28 56.63 54.15 40.67 67.41 55.83 58.26 43.12 64.73 44.18 35.47 24.26
CPPNet [10] 73.28 61.74 62.15 40.67 27.89 18.43 66.47 53.82 51.63 38.27 64.35 52.78 54.86 40.25 62.73 42.68 32.17 21.94
CellViT [2] 75.19 63.85 62.94 41.57 29.73 19.68 69.16 56.42 54.29 40.83 67.58 55.94 62.47 45.16 64.85 44.27 35.62 24.38
PromptNucSeg [1] 74.63 63.17 61.98 40.85 28.94 19.26 68.35 55.73 53.74 40.15 66.82 55.29 63.63 45.48 64.12 43.69 34.86 23.74
PathoSAM [15] 76.54 65.62 60.76 41.10 31.20 20.55 73.13 58.52 67.09 50.33 70.12 58.25 63.86 46.56 63.59 41.30 40.39 27.36
Co-Seg++ 80.25 70.59 64.75 45.03 36.91 23.65 77.06 61.63 69.71 53.66 74.68 61.37 67.81 50.23 69.46 46.53 47.71 33.31
Table 4: Comparison with state-of-the-arts on CBCT instance segmentation.
Methods Lower Teeth Upper Teeth
F1 AJI F1 AJI
nnUnet [7] 74.28 68.15 62.74 56.83
Swin-Umamba [12] 72.85 66.42 60.37 54.29
EMCAD [8] 73.96 67.58 61.84 55.67
H-SAM [3] 75.43 69.15 63.28 57.15
Zig-RiR [39] 74.67 68.29 62.55 56.41
ToothNet [28] 77.16 71.05 64.30 58.32
HMG-Net [41] 76.83 70.67 65.12 59.24
SGA-Net [42] 77.94 72.36 66.47 60.85
T-Mamba [24] 78.26 72.71 67.15 61.38
ToothSC-SAM [36] 79.15 73.42 68.29 62.74
TSG-GCN [30] 78.56 71.41 69.18 62.48
Co-Seg++ 81.99 74.15 72.41 66.92
4 Experiments
4.1 Datasets and Implementations
4.1.1 Datasets
To validate the effectiveness of the proposed Co-Seg++, we conduct comprehensive evaluations across diverse medical domains, including three H&E-stained histopathology datasets: PUMA [16], GlaS [17], CRAG [14] and one CBCT dataset: ToothFairy2 [5], as follows:
PUMA [16] is a melanoma nuclei and tissue segmentation dataset containing 206 pathological slides scanned at 40$\times$ magnification with a resolution of $1024\times 1024$. The dataset covers two tissue regions (i.e., tumor and stroma) and three nuclei types (i.e., tumor, lymphocyte, and other cells).
GlaS [17] dataset, derived from the 2015 Gland Segmentation Challenge, comprises 165 histopathology images captured at 20$\times$ magnification with glandular regions and epithelial, connective and lympho-reticular nuclei instances. The image size varies from $574\times 433$ to $775\times 522$.
CRAG [14] dataset contains 216 colorectal adenocarcinoma patches at 20x magnification with a resolution of $996\times 996$. The dataset annotates glandular regions and epithelial, connective and lympho-reticular nuclei instances.
ToothFairy2 [5] comprises 480 cases with comprehensive annotations for dental anatomical structures, including jawbone and pharynx regions and tooth instances. The image size varies from $272\times 345$ to $512\times 512$.
4.1.2 Implementation Details
We perform all experiments on a NVIDIA A5000 GPU using PyTorch. For fair comparisons, all semantic and instance segmentation methods are implemented with the same training settings and configurations. We utilize the pre-trained SAM ViT-H [32] structure as the image encoder of medical SAM baselines. We adopt the Hirea-L structure [43] as the image encoder of our Co-Seg++. We apply the optimizer using Adam with an initial learning rate of $1\times 10^{-4}$ and use the exponential decay strategy to adjust the learning rate with a factor of 0.98. The loss coefficient $\lambda_{1}$ is set to 0.5. We split all datasets into the training, validation, and test sets as 7:1:2, and all images are resized to $1024\times 1024$ during the fine-tuning and evaluation stage [45]. The batch size and the training epoch are set to 16 and 300. Our Co-Seg++ introduces an additional 8.3% learnable parameters compared to the baseline (in Table 7).
4.1.3 Evaluation Metrics
To perform a comprehensive evaluation of versatile medical segmentation, we apply different metrics in terms of three segmentation tasks. For semantic segmentation, we select the Dice coefficient, mean intersection over union (mIoU), and Hausdorff distance (HD). For instance segmentation, we compare the performance with four metrics, including object-level F1 score and aggregated Jaccard index (AJI). For panoptic segmentation, we adopt panoptic quality (PQ). Except for HD, which measures the distance between predicted and ground truth boundary points, higher scores for these metrics indicate better segmentation quality.
Table 5: Comparison with state-of-the-arts on histopathology panoptic segmentation.
PUMA CRAG GlaS
Methods Semantic Instance Semantic Instance Semantic Instance
Tumor Stroma Tumor Lymph Other Cells Gland Epithelial LR Connective Gland Epithelial LR Connective
nnUnet [7] 57.84 35.72 59.47 49.31 22.64 64.26 52.38 48.79 51.27 55.39 44.28 46.94 29.03
Swin-Umamba [12] 58.96 36.45 60.73 50.68 23.17 65.84 53.71 49.82 52.59 56.85 45.13 47.86 29.67
EMCAD [8] 59.42 36.89 61.15 51.04 23.85 66.38 54.29 50.34 53.18 57.48 45.72 48.41 30.15
H-SAM [3] 60.73 37.52 62.58 51.94 25.94 67.29 55.18 51.26 53.97 58.16 46.42 49.28 30.84
Zig-RiR [39] 60.15 37.18 62.07 51.63 25.31 66.82 54.75 50.89 53.54 57.73 46.09 48.95 30.46
HoverNet [9] 60.85 37.68 62.74 52.15 26.37 67.52 55.47 51.63 54.29 58.47 46.71 49.58 30.96
HookNet [40] 58.73 36.21 60.38 50.42 24.72 65.29 53.18 49.35 52.14 56.28 44.89 47.53 29.48
CellPose [11] 61.29 37.94 63.15 52.48 26.85 68.14 55.83 51.17 54.72 58.96 47.15 50.14 31.18
CPPNet [10] 60.74 37.16 62.47 51.69 26.34 67.45 55.24 52.48 53.16 60.38 45.52 48.47 30.52
CellViT [2] 61.95 38.28 64.72 52.81 27.69 68.73 56.49 52.69 55.41 61.67 47.73 50.68 31.74
PromptNucSeg [1] 61.83 38.19 63.58 52.74 27.52 68.56 56.38 52.59 55.28 62.52 47.64 50.59 31.65
PathoSAM [15] 62.17 38.45 65.01 52.86 27.88 69.98 56.22 52.45 56.05 62.42 47.54 50.42 31.20
Co-Seg++ 64.52 41.67 67.37 54.60 29.48 72.69 59.61 54.90 58.84 65.32 49.01 53.04 34.24
Table 6: Comparison with state-of-the-arts on CBCT panoptic segmentation.
Methods Lower Upper Pharynx
Jawbone Teeth Jawbone Teeth
nnUnet [7] 75.82 61.46 48.73 52.35 84.27
Swin-Umamba [12] 76.94 62.85 49.58 53.72 85.69
EMCAD [8] 77.63 63.49 50.21 54.38 86.45
H-SAM [3] 78.85 64.92 51.47 55.83 87.94
Zig-RiR [39] 78.32 64.17 50.89 55.24 87.38
ToothNet [28] 77.47 62.73 48.14 52.47 85.62
HMG-Net [28] 78.18 64.29 49.86 54.12 86.25
SGA-Net [42] 80.24 66.58 52.74 57.29 89.43
T-Mamba [24] 79.67 66.15 52.18 56.81 88.87
ToothSC-SAM [36] 80.93 67.42 53.04 57.65 90.16
TSG-GCN [30] 81.00 67.74 53.26 57.98 90.31
Co-Seg++ 83.25 70.84 57.57 61.66 94.49
4.2 Comparison on Medical Semantic Segmentation
To comprehensively evaluate the performance of Co-Seg++, we first perform comparisons with state-of-the-art methods on medical semantic segmentation across three histopathology datasets. As shown in Table 1, classic semantic segmentation methods [7, 8] generally underperform compared to prompt-based frameworks [3]. On the PUMA dataset, H-SAM [3] achieves competitive results with 91.31% Dice for tumor segmentation, while specialized histopathology methods such as PromptNucSeg [1] and PathoSAM [15] demonstrate superior performance with 91.57% and 91.78% Dice, respectively. In contrast, our Co-Seg++ substantially outperforms all baseline methods across all datasets. More notably, for the challenging stroma segmentation task, Co-Seg++ demonstrates significant superiority with 60.88% Dice, representing a substantial 5.54% improvement over the best baseline Zig-RiR [39].
Table 2 demonstrates the effectiveness of Co-Seg++ on dental CBCT semantic segmentation. We observe that general medical segmentation methods (e.g, nnUNet [7]) achieve competitive baseline performance, while recent Mamba and prompt approaches further promote dental segmentation. Remarkably, Co-Seg++ achieves state-of-the-art performance across all anatomical structures with statistical significance (e.g., P-value $<$ 0.005). For lower jawbone segmentation, Co-Seg++ reveals 96.54% Dice and 93.85% mIoU, representing a 1.76% Dice improvement over the best baseline ToothSC-SAM [36]. The improvement is even more pronounced for upper jawbone segmentation, where Co-Seg++ achieves 92.78% Dice with a 2.31% improvement. For pharynx segmentation, Co-Seg++ demonstrates consistent superiority with 97.73% Dice and the lowest HD of 13.27, indicating exceptional precision in boundary delineation. These comparisons validate the superior performance of our Co-Seg++ on diverse medical semantic segmentation tasks through mutual optimization.
4.3 Comparison on Medical Instance Segmentation
We further evaluate Co-Seg++ on medical instance segmentation tasks to demonstrate the effectiveness of our collaborative learning framework in handling complex instance-level segmentation challenges. Table 3 presents comprehensive results on histopathology instance segmentation across multiple cell types and datasets. We observe that specialized histopathology methods generally outperform generic segmentation approaches. CellPose [2] and PathoSAM [15] demonstrate superior performance compared to traditional methods such as nnUNet [7]. Our Co-Seg++ consistently surpasses all baselines by substantial margins across all datasets and cell categories with a P-value $<0.001$. On the PUMA dataset, Co-Seg++ achieves remarkable performance for tumor instance segmentation with 80.25% F1 score and 70.59% AJI, representing significant improvements of 3.71% F1 and 4.97% AJI over the best baseline PathoSAM [15].
As shown in Table 4, Co-Seg++ demonstrates superior performance on dental CBCT instance segmentation tasks. We observe progressive improvements from general medical methods to specialized dental segmentation approaches, with ToothSC-SAM [36] and T-Mamba [24] representing strong baselines. Co-Seg++ achieves exceptional results with 81.99% F1 score for lower teeth segmentation, representing improvements of 2.84% F1 score over the best baseline ToothSC-SAM [36]. For upper teeth segmentation, which is typically more challenging due to anatomical complexity, Co-Seg++ demonstrates substantial superiority with 72.41% F1 and 66.92% AJI, achieving improvements of 3.23% F1 and 4.44% AJI, respectively. These results prove the effectiveness of our cross-guidance approach in handling complex instance boundaries and overlapping structures common in medical imaging.
Figure 6: Visualization of tissuegland semantic and nuclei instance segmentation. Our Co-Seg++ exhibits the best results, recognizing more nuclei instances with accurate categories and tissuegland boundaries while having fewer false positives.
Table 7: Ablation Study of Co-Seg++ on Histopathology Panoptic Segmentation.
$P$ $D$ $C$ PUMA CRAG GlaS
Semantic Instance Semantic Instance Semantic Instance
Tumor Stroma Tumor Lymph Other Cells Gland Epithelial LR Connective Gland Epithelial LR Connective
1 61.85 37.92 64.58 52.41 27.35 68.45 55.72 51.88 55.16 61.94 46.85 49.74 31.08
2 62.78 38.65 65.89 53.15 28.12 69.87 56.89 52.94 56.28 63.18 47.69 50.85 32.05
3 62.34 38.28 65.12 52.73 27.68 69.02 56.14 52.25 55.63 62.47 47.21 50.18 31.46
4 63.24 39.18 66.45 53.52 28.59 70.52 57.48 53.67 57.01 63.89 48.15 51.48 32.68
5 63.15 39.02 66.28 53.38 28.47 70.31 57.29 53.44 56.79 63.72 48.02 51.29 32.52
6 63.82 40.12 66.91 53.99 29.08 71.84 58.47 54.32 57.92 64.55 48.69 52.18 33.31
7 63.58 39.54 66.82 53.89 28.92 71.18 57.92 54.21 57.45 64.41 48.58 51.96 33.15
8 64.52 41.67 67.37 54.60 29.48 72.69 59.61 54.90 58.84 65.32 49.01 53.04 34.24
Figure 7: Visualization of CBCT semantic and instance segmentation. Benefiting from the mutual guidance of both segmentation tasks, Co-Seg++ delineates precise jawbone areas and segments accurate teeth.
4.4 Comparison on Medical Panoptic Segmentation
To comprehensively evaluate the versatility of Co-Seg++, we conduct extensive experiments on medical panoptic segmentation, which simultaneously requires both semantic understanding and instance-level discrimination capabilities. Table 5 demonstrates the superior performance of Co-Seg++ on histopathology panoptic segmentation across diverse tissue, gland and nuclei types. We observe that specialized methods like CellViT [2] and PathoSAM [15] achieve competitive performance compared to general segmentation approaches, with PathoSAM [15] showing strong results across most categories. In contrast, Co-Seg++ consistently outperforms all baseline methods with substantial improvements and statistical significance (e.g., with a P-value $<0.001$) across all anatomical structures and datasets. On the PUMA dataset, Co-Seg++ achieves remarkable performance for tumor panoptic quality with 64.52% (2.35% improvement over PathoSAM [15]) and demonstrates particularly significant improvement for the challenging stroma segmentation task with 41.67% (3.22% improvement). Moreover, we compare the quantitative results of our Co-Seg++ with the best baseline methods in Fig. 6 and our UN-SAM can segment nuclei more accurately with better tissue and gland regions.
We further perform comparisons on dental CBCT panoptic segmentation, as shown in Table 6. SGA-Net [42], ToothSC-SAM [36], and TSG-GCN [30] achieve competitive results, with TSG-GCN showing particularly strong performance across multiple anatomical structures. Co-Seg++ achieves state-of-the-art performance across all dental anatomical components, by a remarkable PQ increase of 3.1% and 3.68% on lower and upper teeth segmentation, respectively. We also elaborate on the panoptic segmentation results in Fig 7. The comprehensive evaluation across panoptic segmentation tasks validates the effectiveness of our mutual prompt-guided collaborative learning in handling the complex interplay between semantic regions and instance boundaries.
4.5 Ablation Study
To investigate the effectiveness of the STP-Encoder $P$, MTC-Decoder $D$, and co-segmentation paradigm $C$, we conduct comprehensive ablation studies on histopathology panoptic segmentation across three datasets (PUMA, CRAG, and GlaS), as illustrated in Table 7. By removing the tailored modules from Co-Seg, we construct an original encoder-shared framework with Hiera-L as the ablation baseline. By separately introducing the STP-Encoder ($2^{nd}$ row), MTC-Decoder ($3^{rd}$ row), and co-segmentation paradigm ($4^{th}$ row), the performance demonstrates consistent improvements across all datasets. Notably, our co-segmentation paradigm achieves the most significant individual contribution, with substantial improvements averaging PQ of 1.61%. We additionally investigate the effects of combined components. By comparing $6^{th}$ and $7^{th}$ rows with the $5^{th}$ row, configurations with co-segmentation demonstrate significant gains, achieving an average PQ of 53.48% and 53.20%, respectively. On this basis, our complete Co-Seg++ framework ($8^{th}$ row) simultaneously adopts all three components to achieve the best performance across all datasets and categories. The full model attains optimal performance with an average PQ of 54.25%, representing an average improvement of 3.10% over the baseline. These comprehensive ablation experiments validate that the tailored STP-Encoder, MTC-Decoder, and co-segmentation paradigm collectively contribute to the superior performance of Co-Seg++ by effectively capturing long-range dependencies, enabling cross-task collaboration, and facilitating mutual enhancement between semantic and instance segmentation.
5 Conclusion
In this work, we identify the interdependencies between semantic and instance segmentation in medical imaging, and propose a novel co-segmentation paradigm to establish the Co-Seg++ framework, allowing semantic and instance segmentation tasks to mutually enhance each other. The framework integrates an STP-Encoder to capture long-range spatial and temporal relationships as prior constraints, and an MTC-Decoder that leverages cross-guidance to strengthen contextual consistency across tasks. Extensive experiments on diverse histopathology and CBCT datasets demonstrate that Co-Seg++ surpasses state-of-the-art methods across versatile segmentation, validating the effectiveness of our collaborative learning approach for comprehensive medical image understanding.